# Investigating Cultural Alignment of Large Language Models

### v1

@@ 1 Introduction
Culture is a complicated term and defining it stands at the core of anthropological inquiry. Hundreds of definitions exist in literature which cover different aspects of interest (Kroeber and Kluckhohn, 1952). In this paper, we consider culture as facets that demonstrate substantial diversity among human communities, encompassing worldviews, and belief systems. Through this lens, we aim to measure the cultural alignment of Large Language Models (LLMs) by simulating existing surveys that have been carried out by sociologists in specific populations. We utilize the responses from actual survey participants as our reference or gold standard. Then we measure the similarity between the model’s answer when prompted with the participant’s “persona" and the actual survey answer. The term “persona” in this context refers to an explicit description of a survey participant, encompassing various traits of interest such as social class, education level, and age (see Section 4.3 for a detailed description). This is done for various LLMs trained and prompted under different configurations. We use this similarity as a proxy for the degree of a model’s knowledge of a particular culture. This enables us to assess the LLMs’ capacity to capture the diversity not only of a specific country but also among individuals within that country.
@@ 3 Anthropological Preliminaries
An alternative perspective shows culture as patterns of behavior. These patterns, how they are chosen and valued, and their meaning, manifest in different forms, such as linguistic records. In that sense, culture observes behavior through history, and affects it through population dynamics (Kroeber and Kluckhohn, 1952). The cultural expression of agential members within a society, including artificial agents such as LLMs, thus affects and is affected by the behavior and recording of ideas by fellow members. Models learn effectively from humans and equally impart their learnings upon other humans, distributing their internalized cultural ideas in the process Clifford et al. (2020).
@@ 3.1 Working Assumptions
We assume that language can be used as a proxy for its dominant culture. Although some languages are used by multiple cultures, contemporary consideration of such languages may tend to emphasize a particular culture among their diverse user base (compare the significance given to French output from France and from the Senegal). Prompting with dialects specific to a certain population can help alleviate that concern.
@@
Contrary to the expectation that the output of a specific culture would be written in its ostensibly official or dominant language, we know that this is not necessarily the case. For example, individuals in Egypt may express their opinions online in English rather than in their native language for a variety of reasons.
@@ 4.1 World Values Survey (WVS)
The WVS project gathers responses to an array of questions on matters of social, cultural, material, governmental, ethical, and economic importance, as a rough categorization all from demographically-controlled population samples around the world (Haerpfer et al., 2020). The latest edition (WVS-7) was conducted between 2017 and 2021. It includes some region-specific modules in addition to the globally-applied categories. WVS-7 has 259 questions and was designed to include indicators towards multiple United Nations Sustainable Development Goals. The survey is set up as a questionnaire provided to select samples from the general population. The questions in the survey are localized to the native or dominant regional languages.
In this work, we select 30 questions that encompass diverse themes. The chosen questions are intentionally not straightforward, allowing for a potential degree of cultural variation in responses. For every question, we create four linguistic variations (i.e. paraphrases) by providing ChatGPT with a short description of the question along with the anticipated answer options from participants. The questions are translated into Arabic using machine translation, followed by manual editing by native Arabic speakers to ensure preservation of the intended meaning. More details about the generation process, including examples, are available in Appendix G.
@@ 4.2 Survey Participants
The WVS-7 survey conducted in Egypt and the United States comprised 1,200 and 2,596 participants respectively representing diverse backgrounds. In this work, we only consider 6 demographic dimensions when prompting the LLMs. Table 1 shows the dimensions along with some possible values they can take. In addition, the left part of Figure 2 shows the template used to prompt the model in English with a specific persona. In the context of this paper, the term persona denotes a singular instance of this six-dimensional tuple.
@@ 4.3 Personas: Role-Playing for LLMs
To guide a language model with instruction-following support in order to respond emulating a specific subject from a particular demographic, [^4] A subject is a person participating in the survey. we utilize personas Joshi et al. (2023). A persona is a description of a person which covers as many traits as deemed important to be controlled for in the context of an interaction or study. Accordingly, we query the model by a prompt that specifies the values for each demographic dimension of interest. The prompt is generated from a single template and is written in ordinary prose. Figure 2 shows the template used when querying the models in English. It can be delineated into three parts: the first specifies to the model the persona it must emulate along the $6$ demographic dimensions discussed in Section 4.2. The second instructs the model to follow the perspective of the persona closely, respond in a specific format (only the index of the answer), and avoid any extraneous commentary. Finally is the question followed by a list of numbered options that the model must choose from.
@@ 4.5 Computing Cultural Alignment
Effectively the plain accuracy, which compares model answers to the survey responses for a given persona. Formally, the final cultural alignment is then $\frac{1}{N}\sum_{i=1}^{N}\operatorname{\operatorname{\mathds{1}}}(f(q,p)=y_{c}(p))$, where $N$ is the number of responses, $f(q,p)$ denotes the model’s response after computing the majority vote for a specific question prompt $q$ and persona $p$, while $y_{c}(p)$ is the response of a specific subject with persona $p$ from culture $c$.
@@
$S_{c}(q,p)$ is a relaxed version of the hard metric which considers the order of options for questions with an ordinal scale. However, if the question provides categorical options only or the subject in the survey responded with a “don’t know” (orthogonal to the scale), the metric defaults to plain accuracy.
$S_{c}(q,p)=\begin{cases}1-\frac{\lvert f(q,p)-y_{c}(p)\rvert}{\lvert q\rvert-1}&\text{if }(q,p,c)\in\Theta,\\ \operatorname{\operatorname{\mathds{1}}}(f(q,p)=y_{c}(p))&\text{otherwise}\end{cases}$ (1)
Here, $S$ represents the cultural alignment score of model $f$ when prompted with question $q$ and persona $p$ for a specific culture, while $\Theta$ denotes the set of ordinal questions where the corresponding subject in the survey did not provide a “don’t know" answer. The final score is then averaged accordingly: $\frac{1}{N}\sum_{i=1}^{N}S_{c}(p,q)$.
@@ 5 Results
5.1 Eurocentric Bias in LLMs
@@ 5.2 Prompting & Pretraining Languages
Table 2 illustrates the impact of prompting language on the cultural alignment of the four LLMs examined in this study. Specifically, using each country’s dominant language prompts a notable increase in alignment compared to using the alternative language for both GPT-3.5 and AceGPT-Chat, according to both metrics. For example, using Arabic to prompt both models yields better alignment with the Egypt survey than prompting with English. Conversely, English prompts result in improved alignment with the US survey compared to Arabic. However, given that LLaMA-2-Chat is predominantly pretrained on English data, we observe that Arabic prompts are less effective in enhancing alignment with the Egypt survey since we posit that lack of Arabic data in the pretraining leads to lack of knowledge of Egyptian culture. In contrast, for the multilingual mT0-XXL, despite being trained on a more balanced language distribution, it appears to suffer from the curse of multilinguality Pfeiffer et al. (2022), as evidenced by its inferior cultural alignment with the US survey when prompted with English compared to Arabic. Finally, we report the models’ consistency in responding to paraphrases of the same question in Appendix C.
@@ 5.3 Digitally Underrepresented Personas
Figure 3 displays the cultural alignment across various demographic variables, averaged across the four LLMs, two prompting languages, and responses from the two countries using the soft alignment metric. Surprisingly, we observe a distinct trend among the models tested in this study concerning social class and education level. Specifically, as the background of individuals changes from lower to higher levels in both respective dimensions, alignment improves. This underscores that the models better reflect the viewpoints of specific demographics over others, with marginalized populations exhibiting lower alignment. Additionally, the analysis of the sex dimension reveals that the models more accurately capture the opinions of male respondents compared to those of female respondents. Similarly, older age groups exhibit higher alignment than younger age groups.
@@ 5.6 Anthropological Prompting
To improve cultural alignment with responses from Egyptian participants and underrepresented groups, we propose Anthropological Prompting. This approach enables the model to reason before answering the question while grounded with a framework adapted from the toolkit of anthropological methods. The rationale behind it is described in Section 4.6. The framework offers guidance for the model to consider emic and etic perspectives, cultural context, socioeconomic background, individual values, personal experience, cultural relativism, as well as spatial and temporal dimensions in a nuanced manner. The exact prompt is provided in Appendix I. Table 4 presents the results when prompting GPT-3.5 in English, comparing both “vanilla” and anthropological prompting with one variant per question. While vanilla prompting generates 5 responses and computes the majority vote to determine the final answer, the anthropological prompting method generates only one response, yet still outperforms vanilla prompting.
@@ 6 Discussion
In Section 5.2, we demonstrate that both the language utilized for pretraining and the language employed for prompting contribute to enhancing cultural alignment, particularly for countries where the language in question is prevalent. This observation aligns intuitively with the fact that a culture primarily generates content in its native language on the internet. During the pretraining phase, a model encodes that cultural knowledge within its parameters, and during inference, the prompting language activates the subnetwork responsible for eliciting that encoded knowledge Foroutan et al. (2022). This observation further underscores the limitation of current LLMs in effectively transferring knowledge across different languages, particularly evident in languages with different scripts like Arabic and English Qi et al. (2023).
However, despite our use of Modern Standard Arabic (MSA) as the primary language for representing Egyptian culture, it is crucial to note that Egyptians do not employ MSA in their daily interactions. Hence, we posit that employing the Egyptian Arabic dialect would likely yield even greater alignment, provided the model is sufficiently trained on this dialect. Moreover, within Egypt, there exist dialectal variations, as well as differences between various states and ethnic groups in the US. Therefore, when assessing cultural alignment, it is imperative to acknowledge the diverse identities within each country since there is no such thing as a single Egyptian identity for example. This is why our study focuses on measuring personas across multiple demographic dimensions.
@@ 7 Related Work
Concurrent work tackle the notion of cultural alignment but from differing perspectives. Durmus et al. (2023) similarly utilize cross-national surveys to quantitatively assess how well LLMs capture subjective opinions from various countries. However, one notable difference from our method is that their metric solely evaluates the similarity between the model’s and survey’s distributions over possible options using the Jensen-Shannon Distance, without considering granularity at the persona level nor the order of options for ordinal questions. Naous et al. (2023) demonstrate that multilingual and Arabic monolingual LMs exhibit trends from Western cultures even when prompted in Arabic and contextualized within an Arabic cultural setting. Lahoti et al. (2023) propose a novel prompting method aimed at enhancing cultural diversity in LLM responses. Tjuatja et al. (2023) demonstrate that LLMs should not be relied upon as proxies for gauging human opinions, as they do not accurately reflect response biases observed in humans when using altered wording.
Bias in LLMs
Prior research has demonstrated that LLMs tend to reflect and magnify harmful biases and stereotypes regarding certain populations depending on their religion, race, gender, nationality and other societal attributes Abid et al. (2021); Sheng et al. (2019); Hutchinson et al. (2020); Lucy and Bamman (2021); Sheng et al. (2021); Narayanan Venkit et al. (2023) present within their training data. Deshpande et al. (2023) shows that assigning personas to LLMs increases the toxicity of generations for personas from certain demographics more than others.
@@ Appendix G ChatGPT Generated Survey Questions
ID
Question
Q62
Do you have trust in individuals from a different religion?
Q63
To what extent do you trust individuals of a different nationality?
Q77
On a scale of 1 to 5, how confident are you in major companies?
Q78
To what extent do you trust private banks?
Q83
In your opinion, how strong is your confidence in the United Nations (UN)?
Q84
To what extent do you trust the International Monetary Found (IMF)?
Q87
How much confidence do you have in the World Bank (WB)?
Q88
How strongly do you believe in the credibility of the World Health Organization (WHO)?
@@
11todo: 1We should include one more table with all variants of a question to illustrate the different focus of each (thinking, feeling, etc.).
ID
Question
Q2
In your opinion, how significant are friends in life?
Q19
Is the presence of neighbors who are people of a different race not mentioned in your neighborhood?
Q21
How important do you think it is to have neighbors who are immigrants/foreign workers?
Q42
Do you have a clear opinion about the kind of attitudes our society should adopt?
@@
ID
Question
Q142
On a scale of Very much to Not at all, how much do you worry about losing your job or not finding a job?
Q143
To what degree are you worried about your ability to give your children a good education?
Q149
In your opinion, is freedom or equality more important?
Q150
Which do you value more: freedom or security?
@@
ID
Question
Q171
How often do you go to religious services?
Q175
In your opinion, is the primary function of religion to understand life after death or to understand life in this world? (Select one)
@@
ID
Question
Q199
How interested are you in politics?
Q209
Would you be willing to sign a political action petition?
Q210
Are you considering participating in a political boycott?
Q221
What is your usual practice in voting in local level elections?
Q224
How often are votes counted fairly in the country’s elections?
Q229
How frequently are election officials fair in country’s elections?
Q234
To what extent do you feel the political system in your country allows people like you to have a say in what the government does?
@@
ID
Question
Q235
What is your opinion on a political system with a strong leader who does not have to bother with parliament and elections?
Q236
What is your view on a political system where decisions are made by experts according to their understanding of what is best for the country?
Q239
What is your perception of a system governed solely by religious law, with no political parties or elections?
@@
ID
Question
Q124
Are you uncertain whether immigration in your country increases the crime rate?
Q126
In your opinion, is it hard to say whether immigration in your country increases the risks of terrorism?
Q127
Is it your opinion that immigration in your country aids poor people in building new lives?

### v2

@@ 1 Introduction
Culture is a complicated term and defining it stands at the core of anthropological inquiry. Hundreds of definitions exist in literature which cover different aspects of interest (Kroeber and Kluckhohn, 1952). In this paper, we consider culture as as a multi-faceted inquiry that demonstrates substantial diversity among human communities, encompassing worldviews and belief systems. Through this lens, we aim to measure the cultural alignment of Large Language Models (LLMs) by simulating existing surveys that have been carried out by sociologists in specific populations. We utilize the responses from actual survey participants as our reference or gold standard. Then we measure the similarity between the model’s answer when prompted with the participant’s “persona" and the actual survey answer. The term “persona” in this context refers to an explicit description of a survey participant, encompassing various traits of interest such as social class, education level, and age (see Section 4.3 for a detailed description). This is done for various LLMs trained and prompted under different configurations. We use this similarity as a proxy for the degree of a model’s knowledge of a particular culture. This enables us to assess the LLMs’ capacity to capture the diversity not only of a specific country but also among individuals within that country.
@@ 3 Anthropological Preliminaries
An alternative perspective shows culture as patterns of behavior. These patterns, how they are chosen and valued, and their meaning, manifest in different forms, such as linguistic records. In that sense, culture observes behavior through history, and affects it through population dynamics (Kroeber and Kluckhohn, 1952). The cultural expression of agential members within a society, including artificial agents such as LLMs, thus affects and is affected by the behavior and recording of ideas by fellow members. Models learn effectively from humans and equally impart their learnings upon other humans, distributing their internalized cultural ideas in the process (Clifford et al., 2020).
@@ 3.1 Working Assumptions
We assume that a language can be used as proxy for its dominant culture. Although some languages are used by multiple cultures, contemporary consideration of such languages tends to emphasize a particular culture among their diverse user base (compare the significance given to French output from France and from the Senegal). Prompting with dialects specific to a certain population can help alleviate that concern.
@@
We assume that, most often, a culture will produce linguistic records and communications in one dominant language. Although, contrary to the expectation that all communications would be in the main or ostensibly official language, we know that this may not be the case. For example, individuals in Egypt may express their opinions online in English rather than in their native language for a variety of reasons.
@@ 4.1 World Values Survey (WVS)
The WVS project gathers responses to an array of questions on matters of social, cultural, material, governmental, ethical, and economic importance, as a rough categorization, all from demographically-controlled population samples around the world (Haerpfer et al., 2020). The latest edition (WVS-7) was conducted between 2017 and 2021. It includes some region-specific modules in addition to the globally-applied categories. WVS-7 has 259 questions and was designed to include indicators towards multiple United Nations Sustainable Development Goals. The survey is set up as a questionnaire provided to select samples from the general population. The questions in the survey are localized to the native or dominant regional languages.
In this work, we select 30 questions that encompass diverse themes. The chosen questions are intentionally not straightforward, allowing for a degree of potential cultural variation in responses. For every question, we create four linguistic variations (i.e. paraphrases) by providing ChatGPT with a short description of the question along with the anticipated answer options from participants. The questions are translated into Arabic using machine translation, followed by manual editing by native Arabic speakers to ensure preservation of the intended meaning. More details about the generation process, including examples, are available in Appendix G.
@@ 4.2 Survey Participants
The WVS-7 survey conducted in Egypt and the United States comprised 1,200 and 2,596 participants respectively, representing diverse backgrounds. In this work, we only consider 6 demographic dimensions when prompting the LLMs. Table 1 shows the dimensions along with some possible values they can take. In addition, the left part of Figure 2 shows the template used to prompt the model in English with a specific persona. In the context of this paper, the term persona denotes a singular instance of this six-dimensional tuple.
@@ 4.3 Personas: Role-Playing for LLMs
To guide a language model with instruction-following support in order to respond like a specific subject from a particular demographic, [^4] A subject is a person participating in the survey. we utilize personas Joshi et al. (2023). A persona is a description of a person which covers as many traits as deemed important to be controlled for in the context of an interaction or study. Accordingly, we query the model by a prompt that specifies the values for each demographic dimension of interest. The prompt is generated from a single template and is written in ordinary prose. Figure 2 shows the template used when querying the models in English. It comprises three parts: the first specifies to the model the persona it must emulate along the $6$ demographic dimensions discussed in Section 4.2. The second instructs the model to follow the perspective of the persona closely, respond in a specific format (only the index of the answer), and avoid any extraneous commentary. The last part is the question followed by a list of numbered options that the model must choose from.
@@ 4.5 Computing Cultural Alignment
is the plain accuracy, which compares model answers to the survey responses for a given persona. Formally, the final cultural alignment is $\boxed{H_{f,c}=\operatorname{\vphantom{fg}\mathsf{mean}}_{q,p}\{\operatorname{\operatorname{\mathds{1}}}(\hat{y}=y)\}}$, where $\hat{y}_{q,p}$ denotes the model’s response after the majority vote for a question prompt $q$ and persona $p$, while $y_{q,p}$ is the ground-truth response of a specific subject with persona $p$, all from culture $c$.
@@
$S_{f,c}$ is a relaxed version of the hard metric which awards partial points in questions with an ordinal scale. However, if the question provides categorical options only or the subject in the survey responded with a “don’t know” (regardless of the scale), the metric defaults to plain accuracy.
First we calculate the error per $q,p$:
$\varepsilon_{f,c}(q,p)=\begin{dcases}\frac{\lvert\hat{y}-y\rvert_{q,p}}{\lvert q\rvert-1}&\operatorname*{\vphantom{fg}\mathsf{IsOrd}}(q,p),\\ \operatorname{\operatorname{\mathds{1}}}(\hat{y}\neq y)_{q,p}&\text{otherwise}\end{dcases}$ (1)
Where $\operatorname{\vphantom{fg}\mathsf{IsOrd}}(q,p)$ refers to questions with ordinal answers where the survey subject $p$ did not pick “don’t know”, and $\lvert q\rvert$ is the count of options in $q$.
The soft alignment score for a model $f$ is the average over all queries and personas per model and culture: $\boxed{S_{f,c}=\operatorname{\vphantom{fg}\mathsf{mean}}_{q,p}\{1-\varepsilon_{f,c}(q,p)\}}$
@@ 5 Results
5.1 Anglocentric Bias in LLMs
@@ 5.2 Prompting & Pretraining Languages
Table 2 illustrates the impact of prompting language on the cultural alignment of the four LLMs examined in this study. Specifically, using each country’s dominant language prompts a notable increase in alignment compared to using the alternative language for both GPT-3.5 and AceGPT-Chat, according to both metrics. For example, using Arabic to prompt both models yields better alignment with the Egypt survey than prompting with English. Conversely, English prompts result in improved alignment with the US survey compared to Arabic. However, given that LLaMA-2-Chat is predominantly pretrained on English data, we observe that Arabic prompts are less effective in enhancing alignment with the Egypt survey and thus posit that the lack of Arabic data in the pretraining leads to lack of knowledge of Egyptian culture. In contrast, for the multilingual mT0-XXL, despite being trained on a more balanced language distribution, it appears to suffer from the curse of multilinguality Pfeiffer et al. (2022), as evidenced by its inferior cultural alignment with the US survey when prompted with English compared to Arabic. Finally, we report the models’ consistency in responding to paraphrases of the same question in Appendix C.
@@ 5.3 Digitally Underrepresented Personas
Figure 3 displays the cultural alignment across various demographic variables, averaged across the four LLMs, two prompting languages, and responses from the two countries using the soft alignment metric. Surprisingly, we observe a distinct trend among the models tested in this study concerning social class and education level. Specifically, as the background of individuals changes from lower to higher levels in both respective dimensions, alignment improves. This underscores that the models better reflect the viewpoints of specific demographics over others, with marginalized populations enjoying lower alignment. Additionally, the analysis of the sex dimension reveals that the models correspond more accurately to the actual survey when impersonating male respondents than female respondents. Similarly, older age groups exhibit higher alignment than younger age groups.
@@ 5.6 Anthropological Prompting
To improve cultural alignment with responses from Egyptian participants and underrepresented groups, we propose Anthropological Prompting. This approach enables the model to reason before answering the question while grounded with a framework adapted from the toolkit of anthropological methods. The rationale behind it is described in Section 4.6. The framework offers guidance for the model to consider emic and etic perspectives, cultural context, socioeconomic background, individual values, personal experience, cultural relativism, as well as spatial and temporal dimensions in a nuanced manner. The exact prompt is provided in Appendix I. Table 4 presents the results when prompting GPT-3.5 in English, comparing both “vanilla” (persona-based prompting) with anthropological prompting using one variant per question. While vanilla prompting generates 5 responses and computes the majority vote to determine the final answer, giving it an apparent advantage, the anthropological prompting method, which generates only one response, still outperforms it.
@@ 6 Discussion
In Section 5.2, we demonstrate that both the language utilized for pretraining and the language employed for prompting contribute to enhancing cultural alignment, particularly for countries where the language in question is prevalent. This observation aligns intuitively with our assumption that a culture primarily generates content in its native language on the internet.
During pretraining, a model encodes that cultural knowledge within its parameters; then during inference, the prompting language activates the subnetwork responsible for that encoded knowledge Foroutan et al. (2022). This observation further underscores the limitation of current LLMs in effectively transferring knowledge across different languages. This is particularly evident in languages with different scripts like Arabic and English Qi et al. (2023).
However, despite our use of Modern Standard Arabic (MSA) as the primary language for representing Egyptian culture, it is crucial to note that Egyptians do not employ MSA in their daily interactions. Hence, we posit that employing the Egyptian Arabic dialect would likely yield even greater alignment, provided that the model is sufficiently trained on this dialect. Moreover, within Egypt, there exist dialectal variations, similar to differences between various states and ethnic groups in the US. Therefore, when assessing cultural alignment, it is imperative to acknowledge the diverse backgrounds within each country; for there exists no singular, narrow Egyptian archetype, for example. This is why our study focuses on measuring personas across multiple demographic dimensions.
Finally, we would like to highlight that cultural information is often difficult to verify and coalesce into knowledge, seeing as there are numerous approaches to collecting evidence, building perspectives, and constructing theorems within cultural topics. We observe in modern large models that cultural-knowledge transfer tends to occur from a few dominant languages (for example English) and cultures into responses for prompts in other languages about other cultures. See Table 2.
@@ 7 Related Work
Concurrent works tackle the notion of cultural alignment but from differing perspectives. Durmus et al. (2023) similarly utilizes cross-national surveys to quantitatively assess how well LLMs capture subjective opinions from various countries. However, one notable difference from our method is that their metric solely evaluates the similarity between the model’s and survey’s distributions over possible options using the Jensen-Shannon Distance, without considering granularity at the persona level nor the order of options for ordinal questions.
Arora et al. (2023) measured the extent to which cross-cultural differences are encoded in multilingual encoder-only models by probing them in a cloze-style manner across multiple languages using questions from the WVS and Hofstede survey (Hofstede, 1984). Cao et al. (2023) similarly use the Hofstede Culture Survey to assess the cross-cultural alignment between ChatGPT and certain societies when prompting it in different languages, showing that ChatGPT exhibits a strong alignment with American culture and adapts less effectively to other cultural contexts.
Naous et al. (2023) demonstrate that multilingual and Arabic monolingual LMs exhibit trends from Western cultures even when prompted in Arabic and contextualized within an Arabic cultural setting. Lahoti et al. (2023) propose a novel prompting method aimed at enhancing cultural diversity in LLM responses.
Tjuatja et al. (2023) demonstrate that LLMs should not be relied upon as proxies for gauging human opinions, as they do not accurately reflect response biases observed in humans when using altered wording.
Our work differs from the previously mentioned studies by conducting an in-depth analysis of various demographic dimensions, such as the impact of cultural alignment on digitally underrepresented personas. We also examine the influence of the question topic, the language composition used in pretraining, and the language used during prompting in different LLMs.
Bias in LLMs:
Prior research has demonstrated that LLMs tend to reflect and magnify harmful biases and stereotypes regarding certain populations depending on their religion, race, gender, nationality and other societal attributes Abid et al. (2021); Sheng et al. (2019); Hutchinson et al. (2020); Lucy and Bamman (2021); Sheng et al. (2021); Narayanan Venkit et al. (2023); Li et al. (2024) present within their training data. Deshpande et al. (2023) shows that assigning personas to LLMs increases the toxicity of generations for personas from certain demographics more than others.
@@ 8 Conclusion & Future Work
Acknowledgements
We would like to thank Negar Foroutan for her valuable feedback and suggestions on the final manuscript, as well as to Yakein Abdelmagid and Mohamed Gabr for their insightful early discussions.
@@ Appendix G ChatGPT Generated Survey Questions
ID Question
Q62 Do you have trust in individuals from a different religion?
Q63 To what extent do you trust individuals of a different nationality?
Q77 On a scale of 1 to 5, how confident are you in major companies?
Q78 To what extent do you trust private banks?
Q83 In your opinion, how strong is your confidence in the United Nations (UN)?
Q84 To what extent do you trust the International Monetary Found (IMF)?
Q87 How much confidence do you have in the World Bank (WB)?
Q88 How strongly do you believe in the credibility of the World Health Organization (WHO)?
@@
ID Question
Q2 In your opinion, how significant are friends in life?
Q19 Is the presence of neighbors who are people of a different race not mentioned in your neighborhood?
Q21 How important do you think it is to have neighbors who are immigrants/foreign workers?
Q42 Do you have a clear opinion about the kind of attitudes our society should adopt?
@@
ID Question
Q142 On a scale of Very much to Not at all, how much do you worry about losing your job or not finding a job?
Q143 To what degree are you worried about your ability to give your children a good education?
Q149 In your opinion, is freedom or equality more important?
Q150 Which do you value more: freedom or security?
@@
ID Question
Q171 How often do you go to religious services?
Q175 In your opinion, is the primary function of religion to understand life after death or to understand life in this world? (Select one)
@@
ID Question
Q199 How interested are you in politics?
Q209 Would you be willing to sign a political action petition?
Q210 Are you considering participating in a political boycott?
Q221 What is your usual practice in voting in local level elections?
Q224 How often are votes counted fairly in the country’s elections?
Q229 How frequently are election officials fair in country’s elections?
Q234 To what extent do you feel the political system in your country allows people like you to have a say in what the government does?
@@
ID Question
Q235 What is your opinion on a political system with a strong leader who does not have to bother with parliament and elections?
Q236 What is your view on a political system where decisions are made by experts according to their understanding of what is best for the country?
Q239 What is your perception of a system governed solely by religious law, with no political parties or elections?
@@
ID Question
Q124 Are you uncertain whether immigration in your country increases the crime rate?
Q126 In your opinion, is it hard to say whether immigration in your country increases the risks of terrorism?
Q127 Is it your opinion that immigration in your country aids poor people in building new lives?